OpenStudy (anonymous):

why is the partial derivative of y^1/3 --- 1/3(y^2/3) and not 1/3(y)^-2/3 ??

OpenStudy (anonymous):

i meant and not (1/3)*y^ -2/3

OpenStudy (jamesj):

It's not. \[ \frac{\partial \ }{\partial y} y^{1/3} = \frac{1}{3}y^{-2/3} \]

OpenStudy (phi):

same thing

OpenStudy (jamesj):

Oh, I see. It wasn't clear in your first expression, but there the y is in the denominator of the fraction. Then yes, phi, is right, because \[ y^{-2/3} = \frac{1}{y^{2/3}} \]

OpenStudy (anonymous):

ohhhh. im doing the existence and uniqueness theorem. so i get it. nothing with a negative exponent is continuos at y=0. even if its on the top, its on the bottom.

OpenStudy (jamesj):

Right.